OpenStudy (anonymous):

A curve C is defined by the parametric equations x=t^2-4t+1 and y=t^3 what is the equation of the line tangent to the point (-2, 27)?

OpenStudy (anonymous):

the derivative of t^2-4t+1 = 2t-4. Plugging in -2 gives 2(-2)-4 which is 0. therefore y=mx+c y=0x+c y=27 assuming the point (-2,27) exists

OpenStudy (anonymous):

okay i see thanks

OpenStudy (anonymous):

so the equation would be y=27 ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

thanks!